# Fully homomorphic encryption, outsourced databases, high-performance computing

### v1

@@ 5.2 Homomorphic Addition
Let $\mathbf{c}^{(1)}=\textsf{SynthEnc}(\mathbf{m}_{1})$ and $\mathbf{c}^{(2)}=\textsf{SynthEnc}(\mathbf{m}_{2})$ be two synthesized ciphertexts. Their homomorphic sum is:
$\mathbf{c}^{(1)}+\mathbf{c}^{(2)}=\sum_{i}(m^{(1)}_{i}+m^{(2)}_{i})\cdot\mathbf{c}_{i}+(\mathbf{r}^{(1)}+\mathbf{r}^{(2)}).$
By linearity of encryption and noise:
$\mathsf{Dec}(\mathbf{c}^{(1)}+\mathbf{c}^{(2)})=\mathbf{m}_{1}+\mathbf{m}_{2}+e(x),$
where the new noise $e(x)$ is the sum of the individual noise components from both ciphertexts. Assuming noise growth is additive, we have:
$\|e(x)\|_{\infty}\leq\|e^{(1)}(x)\|_{\infty}+\|e^{(2)}(x)\|_{\infty}.$
Thus, additive homomorphism is preserved and the noise growth is predictable and controlled.
@@ 5.3 Homomorphic Multiplication
Let $\mathbf{c}^{(1)}=\textsf{SynthEnc}(\mathbf{m}_{1})$ and $\mathbf{c}^{(2)}=\textsf{SynthEnc}(\mathbf{m}_{2})$ be synthesized ciphertexts represented in two-component form:
@@
Their homomorphic product $\mathbf{c}^{(1)}\cdot\mathbf{c}^{(2)}$ yields a ciphertext with three components:
@@
To convert this ciphertext back into the standard 2-component form $(c^{\prime}_{0},c^{\prime}_{1})$, a relinearization step is required. Given a relinearization key $rk=\mathsf{Enc}(s^{2})$, we perform:
@@
where $\mathsf{RLK}_{0}$, $\mathsf{RLK}_{1}$ are decomposition-based key-switching functions:
$\mathsf{RLK}_{j}(d_{2})=\text{KeySwitch}(d_{2},rk_{j}),\quad j\in\{0,1\}.$
This yields a valid ciphertext of $\mathbf{m}_{1}\circ\mathbf{m}_{2}$, where $\circ$ denotes slot-wise multiplication (or approximate multiplication in CKKS). The resulting noise grows quadratically:
$\|e_{\text{mult}}(x)\|_{\infty}=\mathcal{O}(\|e^{(1)}(x)\|_{\infty}\cdot\|e^{(2)}(x)\|_{\infty}+\text{RelinNoise}).$
In leveled schemes, such as BFV or BGV, a modulus switching step may be applied post-multiplication to reduce noise. Given modulus chain $q_{L}>q_{L-1}>\dots>q_{0}$, switching from level $\ell$ to $\ell-1$ rescales:
@@
and correspondingly:
$\|e(x)\|_{\infty}\leftarrow\frac{q_{\ell-1}}{q_{\ell}}\cdot\|e(x)\|_{\infty}.$
However, modulus switching also introduces rounding error, and in schemes like CKKS, approximate rescaling impacts decryption accuracy.
Hence, homomorphic multiplication in synthesized ciphertexts is functionally identical to native ciphertext multiplication, but inherits the usual cost in both depth and noise. Correctness is maintained as long as the total noise remains below the decryption threshold:
$\|e_{\text{final}}(x)\|_{\infty}<\frac{q_{\ell}}{2t}.$
@@ 6.1 Security Model
We consider the standard notion of semantic security under chosen-plaintext attack (IND-CPA). Let $\mathsf{Enc}:R_{t}\to R_{q}^{k}$ denote a semantically secure public-key encryption scheme. Let $\mathsf{SynthEnc}_{\mathcal{B}}^{\mathcal{R}}(\mathbf{m})$ denote our proposed synthesized ciphertext construction using precomputed basis $\mathcal{B}$ and randomized noise pool $\mathcal{R}$.
We define the adversary’s goal as distinguishing two synthesized ciphertexts corresponding to plaintext vectors of its choice, drawn from the same message space $\mathbb{Z}_{t}^{d}$. Formally, the IND-CPA experiment proceeds as follows:
@@
Let $\mathcal{A}$ be a PPT adversary. The IND-CPA advantage of $\mathcal{A}$ in distinguishing synthesized ciphertexts is defined by the following game:
@@
    The challenger runs $\mathsf{KeyGen}()\rightarrow(pk,sk)$ and sends $pk$ to $\mathcal{A}$.
@@
    The challenger computes =$\mathcal{B}=\{\mathsf{Enc}_{pk}(e_{i})\}_{i=1}^{d}$ and =$\mathcal{R}=\{\mathsf{Enc}_{pk}(0)\}_{j=1}^{s}$, and gives $(\mathcal{B},\mathcal{R})$ to $\mathcal{A}$.
@@
    $\mathcal{A}$ submits two equal-length plaintext vectors $(\mathbf{m}_{0},\mathbf{m}_{1})\in\mathbb{Z}_{t}^{d}\times\mathbb{Z}_{t}^{d}$.
@@
    The challenger selects $b\leftarrow\{0,1\}$ uniformly at random, and returns $\mathbf{c}^{*}=\mathsf{SynthEnc}_{\mathcal{B}}^{\mathcal{R}}(\mathbf{m}_{b})$.
  5. 5.
@@
We define the adversary’s advantage as:
$\mathsf{Adv}_{\mathsf{SynthEnc}}^{\mathsf{IND{-}CPA}}(\mathcal{A})=\left|\Pr[b^{\prime}=b]-\frac{1}{2}\right|.$
6.2 Assumptions and Reduction Strategy
Our security analysis is based on a standard reduction to the IND-CPA security of the underlying FHE encryption scheme.
We assume that the public-key encryption function $\mathsf{Enc}:R_{t}\to R_{q}^{k}$ is semantically secure under chosen-plaintext attacks. That is, for all PPT adversaries $\mathcal{B}$, consider the standard IND-CPA game:
@@ 6.2 Assumptions and Reduction Strategy
    The challenger runs $\mathsf{KeyGen}()\to(pk,sk)$ and sends $pk$ to $\mathcal{B}$.
@@
    $\mathcal{B}$ submits two messages $m_{0},m_{1}\in R_{t}$.
  3. 3.
    A bit $b\leftarrow\{0,1\}$ is chosen uniformly at random, and the challenger returns $\mathbf{c}^{*}=\mathsf{Enc}_{pk}(m_{b})$.
  4. 4.
    $\mathcal{B}$ outputs a guess $b^{\prime}\in\{0,1\}$.
Define the adversary’s advantage as:
$\mathsf{Adv}^{\mathsf{IND{-}CPA}}_{\mathsf{Enc}}(\mathcal{B})=\left|\Pr[b^{\prime}=b]-\frac{1}{2}\right|.$
We assume that $\mathsf{Adv}^{\mathsf{IND{-}CPA}}_{\mathsf{Enc}}(\mathcal{B})$ is negligible in the security parameter $\lambda$.
Our goal is to show that if any adversary $\mathcal{A}$ has non-negligible advantage in the $\mathsf{SynthEnc}$ IND-CPA game (as defined above), then we can construct an adversary $\mathcal{B}$ that breaks the IND-CPA security of $\mathsf{Enc}$ with non-negligible advantage.
This reduction proceeds via a sequence of hybrid games (detailed in the next section), in which each game-hop simulates one coordinate substitution between two synthesized ciphertexts. Each hop relies on the indistinguishability of encryptions under $\mathsf{Enc}$, and the cumulative advantage is bounded via a union bound over $d$ coordinate transitions.
6.3 Provable Security
We now prove that if the underlying encryption scheme $\mathsf{Enc}$ is IND-CPA secure, then the synthesized encryption scheme $\mathsf{SynthEnc}$ is also IND-CPA secure.
Our strategy is to construct a sequence of hybrid games $\{G_{i}\}_{i=0}^{d+1}$, where $G_{0}$ corresponds to the challenge ciphertext generated using $\mathbf{m}_{0}$, and $G_{d+1}$ corresponds to the ciphertext for $\mathbf{m}_{1}$. In each intermediate game $G_{i}$, the first $i$ coordinates of $\mathbf{m}_{0}$ are replaced with those of $\mathbf{m}_{1}$, while the remaining coordinates remain unchanged. By bounding the distinguishing advantage between successive games, we obtain the final security bound via a standard hybrid argument.
@@ 6.3 Provable Security
This is the real world experiment with challenge ciphertext:
$\mathbf{c}^{*}=\sum_{i=1}^{d}(m_{0})_{i}\cdot\mathbf{c}_{i}+\mathbf{r},\quad\mathbf{r}\leftarrow\mathcal{R}.$
Game $G_{i}$.
In this game, the challenge ciphertext is:
$\mathbf{c}^{*}=\sum_{j=1}^{i}(m_{1})_{j}\cdot\mathbf{c}_{j}+\sum_{j=i+1}^{d}(m_{0})_{j}\cdot\mathbf{c}_{j}+\mathbf{r}.$
That is, the first $i$ slots are taken from $\mathbf{m}_{1}$, and the rest from $\mathbf{m}_{0}$.
@@
This corresponds to the synthesized ciphertext of $\mathbf{m}_{1}$:
$\mathbf{c}^{*}=\sum_{i=1}^{d}(m_{1})_{i}\cdot\mathbf{c}_{i}+\mathbf{r}.$
Note that $G_{0}$ and $G_{d+1}$ correspond exactly to the real challenge cases for $b=0$ and $b=1$, respectively.
@@
For each $i\in\{0,\dots,d-1\}$, let $G_{i}$ and $G_{i+1}$ be defined as above. If the underlying encryption scheme $\mathsf{Enc}$ is IND-CPA secure, then for any PPT adversary $\mathcal{A}$:
@@
for negligible function $\epsilon$ in the security parameter $\lambda$.
@@
Define the distinguishing difference:
$\Delta_{i}:=\Pr[\mathcal{A}\text{ outputs }b^{\prime}=1\mid G_{i}]-\Pr[\mathcal{A}\text{ outputs }b^{\prime}=1\mid G_{i+1}].$
The difference between $G_{i}$ and $G_{i+1}$ lies in a single coefficient $(m_{0})_{i+1}$ vs. $(m_{1})_{i+1}$ in the linear combination:
$\mathbf{c}^{*}=\mathbf{c}_{\text{fixed}}+(m_{b})_{i+1}\cdot\mathbf{c}_{i+1}.$
If $\mathcal{A}$ can distinguish between games $G_{i}$ and $G_{i+1}$ with non-negligible probability, we can construct a reduction adversary $\mathcal{B}$ that breaks the IND-CPA security of the underlying encryption scheme $\mathsf{Enc}$. Specifically, $\mathcal{B}$ receives a challenge ciphertext $\mathbf{c}^{*}$ corresponding to either $m_{0}$ or $m_{1}$ in a single-slot encryption instance. It embeds this challenge into a full synthesized ciphertext by constructing $\mathbf{c}=\mathbf{c}_{\text{fixed}}+m^{*}\cdot\mathbf{c}_{i+1}$, where $\mathbf{c}_{\text{fixed}}$ encodes all other coordinates from the hybrid game. If $\mathcal{A}$ correctly distinguishes whether $m^{*}=(m_{0})_{i+1}$ or $(m_{1})_{i+1}$, then $\mathcal{B}$ breaks the IND-CPA security of $\mathsf{Enc}$.
Hence, $\Delta_{i}$ must be negligible. ∎
We now turn to the main result of this section. Building on the sequence of hybrid games $\{G_{i}\}_{i=0}^{d+1}$ and the indistinguishability lemma established for adjacent hybrids, we apply a telescoping argument over the entire chain. The accumulated advantage across all transitions allows us to bound the distinguishing probability of any adversary in the full $\mathsf{SynthEnc}$-IND-CPA experiment. The result is formalized in the following theorem.
@@
[IND-CPA Security of Synthesized Encryption] Let $\mathsf{Enc}$ be a semantically secure public-key encryption scheme. Then the synthesized encryption scheme $\mathsf{SynthEnc}$, constructed via basis combination and random zero injection, is also IND-CPA secure. In particular, for any PPT adversary $\mathcal{A}$, there exists a negligible function $\varepsilon(\lambda)$ such that:
$\mathsf{Adv}^{\mathsf{IND\textnormal{-}CPA}}_{\mathsf{SynthEnc}}(\mathcal{A})\leq\varepsilon(\lambda).$
@@
Let $\{G_{i}\}_{i=0}^{d+1}$ be the hybrid game sequence constructed in the previous subsection, where $G_{0}$ corresponds to a synthesized encryption of $\mathbf{m}_{0}$ and $G_{d+1}$ to that of $\mathbf{m}_{1}$. Each intermediate game $G_{i}$ partially interpolates between the two vectors by encrypting a message whose first $i$ coordinates are from $\mathbf{m}_{1}$ and the remaining $d-i$ coordinates from $\mathbf{m}_{0}$.
Let $p_{i}$ denote the probability that the adversary $\mathcal{A}$ outputs $1$ in game $G_{i}$, i.e., $p_{i}:=\Pr[\mathcal{A}\text{ outputs }1\mid G_{i}]$. The adversary’s overall advantage in the $\mathsf{SynthEnc}$-IND-CPA experiment is given by the absolute difference $|p_{0}-p_{d+1}|$.
This difference can be expressed as a telescoping sum over adjacent game transitions:
@@
Applying the triangle inequality, this yields the upper bound:
@@
Each pair of games $(G_{i},G_{i+1})$ differ in a single coordinate of the synthesized plaintext vector. Specifically, only the $(i+1)$-th coordinate is switched from $(m_{0})_{i+1}$ to $(m_{1})_{i+1}$, while all other components are held fixed. Thus, each transition corresponds to a one-coordinate hybrid hop.
By the security of the underlying encryption scheme $\mathsf{Enc}$, and in particular the indistinguishability guaranteed by Lemma 6.3, we know that for each $i$, the distinguishing advantage between $G_{i}$ and $G_{i+1}$ is bounded by a negligible function $\epsilon_{i}(\lambda)$, where $\lambda$ is the security parameter. That is,
$|p_{i}-p_{i+1}|\leq\epsilon_{i}(\lambda)\quad\text{ for all }i\in\{0,1,\dots,d\}.$
We now apply the union bound, a standard tool in cryptographic reductions. In its probabilistic form, it states that for any finite collection of distinguishable events, the probability of at least one occurring is bounded by the sum of their individual probabilities. When applied to game-hopping sequences, this translates to: the total distinguishing advantage across a chain of hybrids is at most the sum of the individual hop advantages. Thus,
$|p_{0}-p_{d+1}|\leq\sum_{i=0}^{d}\epsilon_{i}(\lambda).$
To complete the argument, define a function $\varepsilon(\lambda):=\sum_{i=0}^{d}\epsilon_{i}(\lambda)$. Each $\epsilon_{i}(\lambda)$ is negligible by assumption, and $d$ is polynomially bounded in $\lambda$ since it corresponds to the number of plaintext slots. It is a standard closure property of negligible functions that a finite sum of negligible functions, where the number of terms is bounded by a polynomial, remains negligible. Hence, $\varepsilon(\lambda)$ is negligible.
We conclude that the adversary’s advantage satisfies:
$\mathsf{Adv}^{\mathsf{IND\textnormal{-}CPA}}_{\mathsf{SynthEnc}}(\mathcal{A})=|p_{0}-p_{d+1}|\leq\varepsilon(\lambda),$
as required. ∎

### v2

@@ 5.2 Homomorphic Addition
Let $\mathbf{c}^{(1)}$ and $\mathbf{c}^{(2)}$ be two synthesized ciphertexts:
$\mathbf{c}^{(1)}=\sum_{i=1}^{d}m_{i}^{(1)}\cdot\mathbf{c}_{i}+a_{1}\cdot\mathbf{r}_{0},\qquad\mathbf{c}^{(2)}=\sum_{i=1}^{d}m_{i}^{(2)}\cdot\mathbf{c}_{i}+a_{2}\cdot\mathbf{r}_{0}.$
Their homomorphic sum is:
$\mathbf{c}^{(+)}=\mathbf{c}^{(1)}+\mathbf{c}^{(2)}=\sum_{i=1}^{d}(m_{i}^{(1)}+m_{i}^{(2)})\cdot\mathbf{c}_{i}+(a_{1}+a_{2})\cdot\mathbf{r}_{0}.$
The decryption error grows linearly:
$e_{+}(x)=\sum_{i=1}^{d}(m_{i}^{(1)}+m_{i}^{(2)})\cdot e_{i}(x)+(a_{1}+a_{2})\cdot e_{r}(x),$
yielding a noise bound:
$\|e_{+}(x)\|_{\infty}\leq\sum_{i=1}^{d}|m_{i}^{(1)}+m_{i}^{(2)}|\cdot\|e_{i}(x)\|_{\infty}+|a_{1}+a_{2}|\cdot\|e_{r}(x)\|_{\infty}.$
Correctness is preserved as long as:
$\|e_{+}(x)\|_{\infty}<\frac{q}{2t}.$
@@ 5.3 Homomorphic Multiplication
Let $\mathbf{c}^{(1)}=\textsf{SynthEnc}(\mathbf{m}_{1})$ and $\mathbf{c}^{(2)}=\textsf{SynthEnc}(\mathbf{m}_{2})$ be synthesized ciphertexts, each expressed in standard two-component form:
@@
Their homomorphic product yields a ciphertext in three-component form:
@@
This multiplication introduces quadratic noise growth in $d_{2}$, which must be reduced via a relinearization step to preserve correctness.
Relinearization.
Given a relinearization key $rk=\mathsf{RLK}(s^{2})$, we convert $(d_{0},d_{1},d_{2})$ back to a two-component ciphertext:
@@
where $\mathsf{RLK}_{j}(d_{2})=\mathsf{KeySwitch}(d_{2},rk_{j})$ applies decomposition-based key switching on the high-degree term. This process introduces additional noise, denoted $\eta_{\mathrm{relin}}$, which depends on:
  • the decomposition base $w$,
  • the number of digits $\ell=\lceil\log_{w}q\rceil$,
  • and the key switching noise per digit.
In general, the post-relinearization noise is:
$\|e_{\mathrm{rel}}(x)\|_{\infty}=\mathcal{O}(\|e_{1}\|_{\infty}\cdot\|e_{2}\|_{\infty}+\|\mathbf{m}_{1}\|_{\infty}\cdot\|e_{2}\|_{\infty}+\|\mathbf{m}_{2}\|_{\infty}\cdot\|e_{1}\|_{\infty}+\eta_{\mathrm{relin}}).$
Modulus Switching.
After relinearization, the ciphertext resides at level $\ell$ of the modulus chain $\{q_{L},\dots,q_{0}\}$. To reduce noise, we may perform modulus switching:
@@
which rescales both ciphertext components and their associated noise:
$\|e(x)\|_{\infty}\leftarrow\left(\frac{q_{\ell-1}}{q_{\ell}}\right)\cdot\|e(x)\|_{\infty}+\delta_{\mathrm{round}},$
where $\delta_{\mathrm{round}}$ accounts for rounding errors introduced during scaling. For plaintext modulus $t$ and ciphertext modulus $q_{\ell}$, correctness is preserved if:
$\|e(x)\|_{\infty}<\frac{q_{\ell}}{2t}.$
Homomorphic multiplication in synthesized ciphertexts induces significant noise amplification, especially from $d_{2}$ and its relinearization. Our compile-time ciphertext construction remains fully compatible with standard relinearization and modulus switching procedures. However, the correctness criterion after each multiplication must explicitly account for:
$\|e_{\mathrm{mult}}(x)\|_{\infty}<\frac{q_{\ell}}{2t},$
where $e_{\mathrm{mult}}(x)$ aggregates both multiplicative and key switching noise components. Depth-aware parameter selection remains critical to ensure all ciphertexts produced by $\mathsf{SynthEnc}$ remain decryptable after multiple homomorphic operations.
@@ 6.1 Security Model
We consider the standard IND-CPA (indistinguishability under chosen-plaintext attacks) security model for public-key encryption. Let $\mathsf{Enc}:\mathbb{Z}_{t}^{d}\to R_{q}^{k}$ be a semantically secure encryption scheme, and let $\mathsf{SynthEnc}_{\mathcal{B}}^{\mathbf{r}_{0}}$ denote our ciphertext synthesis procedure using cached basis ciphertexts =$\mathcal{B}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{d}\}$ and a fixed encryption of zero $\mathbf{r}_{0}\leftarrow\mathsf{Enc}(0^{d})$.
The adversary is given $\mathcal{B}$ and $\mathbf{r}_{0}$ as public inputs, and its goal is to distinguish synthesized ciphertexts of chosen plaintexts. The IND-CPA game proceeds as follows:
@@
Let $\mathcal{A}$ be a probabilistic polynomial-time (PPT) adversary. Define the IND-CPA game as:
@@
    The challenger generates a keypair $(pk,sk)\leftarrow\mathsf{KeyGen}()$ and computes basis ciphertexts $\mathbf{c}_{i}\leftarrow\mathsf{Enc}_{pk}(e_{i})$ for $i=1$ to $d$, as well as $\mathbf{r}_{0}\leftarrow\mathsf{Enc}_{pk}(0^{d})$ with fresh randomness. The tuple $(\mathcal{B},\mathbf{r}_{0})$ is sent to $\mathcal{A}$.
@@
    $\mathcal{A}$ submits two plaintext vectors $\mathbf{m}_{0},\mathbf{m}_{1}\in\mathbb{Z}_{t}^{d}$.
@@
    The challenger selects $b\leftarrow\{0,1\}$ and a random scalar $\alpha\leftarrow\mathbb{Z}_{t}$, then returns:
    $\mathbf{c}^{*}=\sum_{i=1}^{d}(m_{b})_{i}\cdot\mathbf{c}_{i}+\alpha\cdot\mathbf{r}_{0}.$
@@
We define the advantage of $\mathcal{A}$ as:
$\mathsf{Adv}_{\mathsf{SynthEnc}}^{\mathsf{IND\textnormal{-}CPA}}(\mathcal{A})=\left|\Pr[b^{\prime}=b]-\frac{1}{2}\right|.$
6.2 Reduction to Underlying Encryption
We now prove that $\mathsf{SynthEnc}$ is IND-CPA secure assuming the underlying encryption scheme $\mathsf{Enc}$ is IND-CPA secure. The proof proceeds in two steps:
@@ 6.2 Reduction to Underlying Encryption
    Show that replacing basis ciphertexts one coordinate at a time leads to negligible change in advantage.
@@
    Show that masking with $\alpha\cdot\mathbf{r}_{0}$ is computationally indistinguishable from a fresh encryption of zero.
We define a sequence of hybrid games $\{G_{i}\}_{i=0}^{d+1}$ and prove a telescoping indistinguishability argument. Let $\mathbf{c}_{i}=\mathsf{Enc}(e_{i})$, and let $\alpha\in\mathbb{Z}_{t}$ be a random scalar.
@@
The challenge ciphertext is:
$\mathbf{c}^{*}=\sum_{i=1}^{d}(m_{0})_{i}\cdot\mathbf{c}_{i}+\alpha\cdot\mathbf{r}_{0}.$
Game $G_{i}$ for $1\leq i\leq d$.
The first $i$ coordinates of $\mathbf{m}_{0}$ are replaced with those of $\mathbf{m}_{1}$:
$\mathbf{c}^{*}=\sum_{j=1}^{i}(m_{1})_{j}\cdot\mathbf{c}_{j}+\sum_{j=i+1}^{d}(m_{0})_{j}\cdot\mathbf{c}_{j}+\alpha\cdot\mathbf{r}_{0}.$
@@
The entire message is switched:
$\mathbf{c}^{*}=\sum_{i=1}^{d}(m_{1})_{i}\cdot\mathbf{c}_{i}+\alpha\cdot\mathbf{r}_{0}.$
@@
If $\mathsf{Enc}$ is IND-CPA secure, then for each $i\in\{0,\dots,d-1\}$:
@@
where $\varepsilon_{1}(\lambda)$ is negligible in the security parameter $\lambda$.
@@
Suppose for contradiction that $\mathcal{A}$ can distinguish $G_{i}$ from $G_{i+1}$ with non-negligible probability. We construct an adversary $\mathcal{B}$ that breaks the IND-CPA security of $\mathsf{Enc}$.
Given challenge ciphertext $\mathbf{c}^{*}\leftarrow\mathsf{Enc}(m^{*})$ for $m^{*}\in\{(m_{0})_{i+1},(m_{1})_{i+1}\}$, $\mathcal{B}$ simulates all other terms in $\mathbf{c}^{*}$:
$\mathbf{c}=\sum_{j=1}^{i}(m_{1})_{j}\cdot\mathbf{c}_{j}+\sum_{j=i+2}^{d}(m_{0})_{j}\cdot\mathbf{c}_{j}+\mathbf{c}^{*}+\alpha\cdot\mathbf{r}_{0}.$
Then runs $\mathcal{A}$ on $(\mathcal{B},\mathbf{r}_{0},\mathbf{c})$. If $\mathcal{A}$ guesses $b^{\prime}=1$ with advantage $\delta$, then $\mathcal{B}$ distinguishes $\mathsf{Enc}(m_{0})$ from $\mathsf{Enc}(m_{1})$ with advantage $\delta$. ∎
[…]
Let $\mathbf{r}_{0}\leftarrow\mathsf{Enc}(0^{d})$ be a ciphertext of zero with fresh randomness. Then for uniformly random $\alpha\leftarrow\mathbb{Z}_{t}$, the product $\alpha\cdot\mathbf{r}_{0}$ is computationally indistinguishable from a fresh encryption of zero:
$\alpha\cdot\mathbf{r}_{0}\approx_{c}\mathsf{Enc}(0^{d}).$
Proof.
Let $\mathbf{r}_{0}=(a,b)=\mathsf{Enc}_{pk}(0^{d})$. Since homomorphic multiplication by a plaintext scalar $\alpha$ is a supported operation, the resulting ciphertext $\alpha\cdot\mathbf{r}_{0}$ is a valid encryption of $0^{d}$ with modified noise.
Let us consider the advantage of an adversary in distinguishing $\alpha\cdot\mathbf{r}_{0}$ from $\mathsf{Enc}(0^{d})$. Suppose this advantage is non-negligible. Then, using the homomorphic scalar multiplication algorithm, one can define an efficient transformation $\mathsf{M}_{\alpha}$ such that:
$\mathsf{M}_{\alpha}(\mathsf{Enc}(0^{d}))=\alpha\cdot\mathsf{Enc}(0^{d}).$
Therefore, the ability to distinguish $\alpha\cdot\mathsf{Enc}(0^{d})$ from fresh encryption contradicts the ciphertext distribution indistinguishability under operations supported by the scheme. Since scalar multiplication does not expose plaintext content or noise structure beyond semantic security guarantees, the output remains computationally indistinguishable from a fresh encryption. ∎
6.3 Main Theorem
@@
[IND-CPA Security of $\mathsf{SynthEnc}$] Let $\mathsf{Enc}$ be a public-key encryption scheme satisfying IND-CPA security. Then the synthesized encryption scheme $\mathsf{SynthEnc}$, constructed via precomputed basis combination and scalar-masked encryption of zero, is also IND-CPA secure. Specifically, for every PPT adversary $\mathcal{A}$:
$\mathsf{Adv}_{\mathsf{SynthEnc}}^{\mathsf{IND\textnormal{-}CPA}}(\mathcal{A})\leq d\cdot\varepsilon_{1}(\lambda)+\varepsilon_{2}(\lambda),$
where $\varepsilon_{1},\varepsilon_{2}$ are negligible functions in the security parameter $\lambda$.
@@
Let $\{G_{i}\}_{i=0}^{d+1}$ be the sequence of hybrid games defined above, where $G_{0}$ corresponds to a synthesized encryption of $\mathbf{m}_{0}$ and $G_{d+1}$ to that of $\mathbf{m}_{1}$. Let $p_{i}:=\Pr[\mathcal{A}\text{ outputs }1\mid G_{i}]$.
The adversary’s distinguishing advantage is given by:
$\mathsf{Adv}_{\mathsf{SynthEnc}}^{\mathsf{IND\textnormal{-}CPA}}(\mathcal{A})=|p_{0}-p_{d+1}|.$
We first apply the telescoping identity:
@@
By the triangle inequality, we obtain:
@@
By Lemma 6.2, each adjacent pair of games $G_{i}$, $G_{i+1}$ differ in exactly one coordinate of the plaintext vector. Since each coordinate term is multiplied by a precomputed ciphertext $\mathbf{c}_{i}=\mathsf{Enc}(e_{i})$, the adversary’s ability to distinguish $G_{i}$ and $G_{i+1}$ is bounded by the IND-CPA security of the underlying encryption scheme $\mathsf{Enc}$. Thus, for each $i$, we have:
$|p_{i}-p_{i+1}|\leq\varepsilon_{1}(\lambda),$
for negligible $\varepsilon_{1}$.
We now apply the union bound over all $d+1$ hybrids:
$|p_{0}-p_{d+1}|\leq\sum_{i=0}^{d}\varepsilon_{1}(\lambda)=(d+1)\cdot\varepsilon_{1}(\lambda).$
Finally, we incorporate the randomized noise term $\alpha\cdot\mathbf{r}_{0}$ used in the synthesized ciphertext. By Lemma 6.2, this term is computationally indistinguishable from a fresh encryption of zero. Therefore, any additional advantage the adversary gains from the masking step is bounded by a negligible function $\varepsilon_{2}(\lambda)$.
Combining both sources of error, we conclude:
$\mathsf{Adv}_{\mathsf{SynthEnc}}^{\mathsf{IND\textnormal{-}CPA}}(\mathcal{A})\leq(d+1)\cdot\varepsilon_{1}(\lambda)+\varepsilon_{2}(\lambda),$
which is negligible in the security parameter $\lambda$, as both $\varepsilon_{1}$ and $\varepsilon_{2}$ are negligible functions. ∎